OpenStudy (anonymous):

solve for x. (ln x)/(ln 3) = lnx-ln3. use a calculatorto evaluate your result, round off to nearest integer

OpenStudy (dumbcow):

using log properties: ln(x) - ln(3) = ln(x/3)

OpenStudy (dumbcow):

then \[\frac{\ln x}{\ln 3} = \ln(x/3)\] take e raised to both sides \[x^{1/\ln 3} = \frac{x}{3}\]

OpenStudy (anonymous):

how would we find the numerical answer for this?

OpenStudy (dumbcow):

divide by x \[x^{1/\ln 3 - 1} = \frac{1}{3}\] \[x^{\frac{1-\ln 3}{\ln 3}} = \frac{1}{3}\] \[x = (\frac{1}{3})^{\frac{\ln 3}{1-\ln 3}}\]